### v1

# LightEndoStereo: A Real-time Lightweight Stereo Matching Method for Endoscopy Images

Abstract
Real-time acquisition of accurate depth of scene is essential for automated robotic minimally invasive surgery, and stereo matching with binocular endoscopy can generate such depth. However, existing algorithms struggle with ambiguous tissue boundaries and real-time performance in prevalent high-resolution endoscopic scenes. We propose LightEndoStereo, a lightweight real-time stereo matching method for endoscopic images. We introduce a 3D Mamba Coordinate Attention module to streamline the cost aggregation process by generating position-sensitive attention maps and capturing long-range dependencies across spatial dimensions using the Mamba block. Additionally, we introduce a High-Frequency Disparity Optimization module to refine disparity estimates at tissue boundaries by enhancing high-frequency information in the wavelet domain. Our method is evaluated on the SCARED and SERV-CT datasets, achieving state-of-the-art matching accuracy and a real-time inference speed of 42 FPS. The code is available at https://github.com/Sonne-Ding/LightEndoStereo.
I Introduction
Minimally invasive surgery (MIS) has become a preferred surgical approach due to its reduced invasiveness and faster recovery times [1]. Endoscopy can provide essential visual guidance in MIS. However, endoscopy faces inherent limitations, including a restricted field of view, lack of tactile feedback, and diminished spatial awareness [2, 3, 4]. To address these challenges, computer-assisted intervention (CAI) techniques have been developed to extract spatial depth information from endoscopic images, with depth estimation emerging as a key focus [3].
Endoscopic stereo matching is the method to obtain tissue depth information from binocular images [5, 6]. Specifically, it generates a disparity map that can be converted into a depth map through a simple mapping relationship. Although there have been numerous studies on stereo matching in the domain of natural images, endoscopic stereo matching still faces some challenges due to its unique application scenarios. Tissue boundaries in endoscopic images are blurred due to smooth transitions [7], making accurate depth estimation difficult. Meanwhile, the increasing adoption of high-definition endoscopes adds challenges for stereo matching to maintain computational efficiency and real-time performance. Therefore, developing an efficient, accurate, and robust endoscopic stereo matching method is essential.
Figure 1: Overview of LightEndoStereo. The figure presents the average MAE and single-frame inference speed of various methods on the SCARED dataset. Our method achieves state-of-the-art (SOTA) matching accuracy while maintaining high inference speed.
Since Jure Zbontar [8] first introduced convolutional neural networks (CNNs) to stereo matching, numerous deep learning-based methods have been proposed in this field. Compared to traditional methods, deep learning approaches, leveraging complex network architectures, adaptive feature extraction, and refined optimization capabilities, have achieved significantly higher matching accuracy [9, 10, 11, 12, 13].Currently, there are two mainstream approaches in stereo matching: those based on 3D convolutions [5, 8, 14, 15, 16, 9, 17, 18] and those based on iterative optimization [19, 20, 21, 22]. Methods based on 3D convolutions can effectively encode essential geometric information for matching. However, they also introduce increased computational complexity [23]. On the other hand, iterative optimization frameworks employ recurrent neural networks (RNNs) to iteratively refine disparity estimation [20], balancing matching accuracy and computational performance by adjusting the number of iterations.When processing high-resolution endoscopic images (e.g., $1024\times 1280$ in the SCARED dataset [24]), these iterative methods often require multiple iterations to achieve satisfactory results. Due to the lack of GPU acceleration, their inference speed lags behind convolution-based methods and fails to meet real-time requirements.
[…]
Our main contributions are summarized as follows:
  • We explore stereo matching methods for real-time depth information acquisition via binocular endoscopy in automated minimally invasive surgery, and propose LightEndoStereo, a lightweight real-time stereo matching algorithm that enables accurate disparity estimation, particularly in high-frequency regions such as tissue boundaries and surgical instrument edges.
  • To achieve high-quality aggregation of matching costs in 3D space with minimal computational overhead, we propose the 3D Mamba Coordinate Attention (MCA) module. This module incorporates spatial information into channel attention and efficiently captures long-range correlations across the entire 3D space.
  • To minimize disparity estimation errors at boundaries, which are particularly evident in endoscopic images, we propose a High-Frequency Disparity Optimization (HFDO) module. This module enhances high-frequency components of context features in the wavelet domain to improve disparity estimation at tissue and instrument boundaries.
II RELATED WORK
II-A Real-time Stereo Matching
For of real-time stereo matching, there are primarily two categories: CNN-based cost aggregation methods and iterative optimization-based methods The use of CNNs in stereo matching traces back to the nascent stages of deep learning. These methods predominantly utilize CNNs to extract features and construct a cost volume for disparity estimation. CNN-based methods can be divided into 2D and 3D architectures contingent upon the processing of the cost volume. For instance, 2D architectures like DispNet [27] employ 2D convolutions to process the cost volume, thereby achieving real-time performance albeit with relatively lower accuracy. Conversely, 3D architectures, such as [28, 29, 16], utilize 3D convolutions to explicitly encode geometric information, which enhances accuracy but at the expense of high computational complexity.
Iterative optimization-based methods, inspired by iterative refinement techniques in optical flow estimation such as RAFT [30], refine disparity estimates through multiple iterations. RAFT-Stereo [21] is a pioneering work in this domain. RAFT employs a GRU (Gated Recurrent Unit) structure to iteratively optimize the disparity map, achieving a balance between high accuracy and efficiency. Building on this foundation, subsequent methods have further enhanced the iterative optimization process. The CREStereo [22] adopts a hierarchical network design, updating the disparity map through a coarse-to-fine recursive approach to better recover complex image details. The IGEV-Stereo [20] introduces adaptive correlation layers and geometry encoding volumes to improve the robustness and accuracy of the models. Similarly, DLNR [19] employs an LSTM (Long-Short-Term Memory) structure to specifically optimize disparity estimates in high-frequency regions, thereby further bolstering the overall performance of the models.
II-B High-Frequency Optimization in Stereo Matching
Several works have focused on optimizing stereo matching for high-frequency components [19, 31, 32].The DLNR [19] highlights that the coupling between the update matrix and the hidden state transfer in the GRU module used by RAFT can lead to inaccuracies in high-frequency detail prediction. To address this, DLNR introduced an LSTM (Long-Short-Term Memory) structure to replace the GRU, thereby achieving better high-frequency disparity prediction performance. The Selective-Stereo [31] highlighted that GRUs with fixed receptive fields struggle to capture both high-frequency information in edges and low-frequency information in smooth regions simultaneously. To overcome this limitation, Selective-Stereo proposed the Selective Recurrent Unit (SRU), which incorporates multi-scale receptive fields into the GRU design. This approach enables the model to optimize across multiple frequency bands, improving the overall accuracy of disparity prediction.
Figure 2: (a) Architecture of LightEndoStereo. The Feature Net is a UNet-like structure with an encoder branch based on MobileNetV4. The group-wise correlation cost volume is utilized. (b) Mamba Coordinate Attention module. Attention is extracted separately across the $H$, $W$, and $D$ dimensions, concatenated along the channel dimension, and then passed through a Bi-Mamba2 layer. Bi-Mamba2 layer extracts dependencies between features in both forward and backward directions across all spatial dimensions ($H+W+D$). The resulting position-sensitive attention maps are used to re-weight the feature map. (c) Wavelet Transform Refine module. The context feature map is decomposed into low-frequency (LL) and high-frequency (LH, HL, HH) components. The LL components are attenuated by a parameter $w$.
III METHOD
In this section, we present the methodology of LightEndoStereo, a lightweight stereo matching algorithm designed for high-precision disparity estimation while maintaining efficient computational performance. We integrate a lightweight feature extraction network, an MCA module for optimized cost aggregation, and an HFDO module to refine disparities in high-frequency regions. The overall architecture is shown in Fig. 2.
III-A Feature Extraction
Multi-scale information is essential for stereo matching. Shallow features capture rich texture and geometric details, while deep features provide important abstract semantic information. As so, we propose a lightweight multi-scale feature extraction network based on the U-Net architecture. The encoder branch utilizes MobileNetV4, an efficient lightweight neural network optimized for mobile devices. In the decoder stage, we employ 2D convolutions and skip connections between adjacent scales to progressively restore multi-scale information.
Given a stereo pair $I_{l},I_{r}\in\mathbb{R}^{3\times H\times W}$, we utilize the MobileNetV4 structure [25] as the downsampling branch of the U-Net multi-scale feature extraction framework. In the upsampling branch, features at scales of $1/16$, $1/8$, and $1/4$ are fused, resulting in the final output features at a $1/4$ scale of the original image, $F_{l},F_{r}\in\mathbb{R}^{C\times H/4\times W/4}$. The feature $F_{l}$ extracted from the encoder branch serves as the context feature for subsequent disparity optimization.
III-B Cost Volume Computation
We employ a Group-wise Correlation cost volume [16] to compute the matching cost between the extracted paired features. For the feature maps $F_{l}$ and $F_{r}$, we divide the channels into $g_{n}$ groups, where $g_{n}=16$ in this work. Within each group, the feature vectors from the left and right images are correlated via the inner product. The cost volume is then constructed by concatenating the correlation results across groups. The cost volume is defined as:
$CV(g,d,x,y)=\frac{1}{C/g_{n}}\langle F_{l}^{g}(x,y),F_{r}^{g}(x-d,y)\rangle,$ (1)
[…]
III-C 3D Mamba Coordinate Attention Guided Cost Aggregation
We propose a novel 3D Mamba Coordinate Attention module. This module extends the coordinate attention framework by embedding positional information into channel attention. Moreover, to capture long-range interactions among channel features across the disparity, width, and height dimensions, we introduce a bidirectional Mamba2 block.
Specifically, for the cost volume $CV\in\mathbb{R}^{g_{n}\times D\times H\times W}$, we treat the disparity dimension $D$ as an additional spatial dimension. We employ three pooling kernels $(H,W)$, $(D,H)$, and $(D,W)$ to perform channel encoding along the horizontal, vertical, and disparity directions, respectively. For a given 3D feature $\mathcal{F}\in\mathbb{R}^{C\times X\times Y\times Z}$ (where $Z$ dimension corresponds to the disparity $D$ in this work), the channel-wise attention extraction for the $c$-th channel along the three directions is formulated as:
$\mathcal{Z}^{x}_{c}(x)$ $=\frac{1}{Y\cdot Z}\max_{y,z}\mathcal{F}(c,x,y,z),$ (2)
$\mathcal{Z}^{y}_{c}(y)$ $=\frac{1}{X\cdot Z}\max_{x,z}\mathcal{F}(c,x,y,z),$
$\mathcal{Z}^{z}_{c}(z)$ $=\frac{1}{X\cdot Y}\max_{x,y}\mathcal{F}(c,x,y,z).$
Subsequently, we introduce non-linearity via a sigmoid function and employ the Mamba2 module to capture long-range dependencies across all spatial dimensions. The refined spatial attention is computed as:
[…]
$A_{x},A_{y},A_{z}=\text{split}(A_{\text{refined}}),$
where $\delta(\cdot)$ denotes the sigmoid function. Finally, the refined attention maps $A_{x}$, $A_{y}$, and $A_{z}$ are applied to the original feature map $\mathcal{F}$ to obtain the optimized feature map:
$\mathcal{F}_{\text{refined}}=\mathcal{F}\odot A_{x}\odot A_{y}\odot A_{z},$ (4)
where $\odot$ denotes broadcast $A_{*}$ and then performs element-wise multiplication. The proposed 3D MCA module is embedded into a light 3 scale 3D UNet cost aggregation network to enhance the model’s ability to capture long-range dependencies across the entire spatial domain. We refer to this aggregation network as 3D UNet-MCA in the following text.
III-D Disparity Computation
To achieve differentiable disparity computation and generate sub-pixel disparity estimates with higher precision, we utilize the soft argmax operation to calculate the disparity [28]. Specifically, the matching cost volume $CV$ is first processed by reducing its feature dimension to 1. Subsequently, the spatial scale is restored through bilinear interpolation, resulting in the final cost volume $CV\in\mathbb{R}^{D\times H\times W}$.
The operations are formulated as follows:
$CV=\text{Interpolate}(\text{Conv}(CV)),$ (5)
where Conv denotes the convolution operation that compresses the feature dimension, and Interpolate represents the bilinear interpolation for spatial scale recovery.
The disparity at each pixel location $(h,w)$ is then computed using the soft argmax operation:
$d(h,w)=\sum_{i=0}^{D-1}i\times\sigma(CV(i,h,w)),$ (6)
where $\sigma$ is the softmax function.
III-E High Frequency Disparity Optimization
In endoscopic images, depth variations are relatively gradual in most regions. This characteristic poses challenges for the model in predicting depth in high-frequency regions, while the boundaries and blood vessels are not prominent in the images. To enhance the optimization of high-frequency regions in the disparity optimization process, while avoiding introducing noise to the well-predicted low-frequency regions, we employ a wavelet module. This module decomposes the context feature $F_{s}$ obtained from $F_{l}$ into low-frequency components(LL) and high-frequency components in horizontal(LH), vertical(HL), and diagonal directions(HH). The low-frequency components are then attenuated by a scale $\omega$, and the inverse wavelet transform is applied to obtain the refined context feature. The refined context feature specifically targets the optimization of high-frequency detail regions in the disparity map.
The context features are derived as follows:
$F_{s}=\text{ReLU}(\text{Linear}(F_{l})),$ (7)
where Conv denotes a 2D $3\times 3$ convolution. The 2D Haar wavelet convolution kernels used are [33]:
$h_{LL}=\frac{1}{2}\begin{bmatrix}1&1\\ 1&1\end{bmatrix},\quad h_{LH}=\frac{1}{2}\begin{bmatrix}1&-1\\ 1&-1\end{bmatrix},$ (8)
$h_{HL}=\frac{1}{2}\begin{bmatrix}1&1\\ -1&-1\end{bmatrix},h_{HH}=\frac{1}{2}\begin{bmatrix}1&-1\\ -1&1\end{bmatrix}.$
The decomposition of the feature map $F_{s}$ is performed as follows:
$F_{s}^{ll}=\omega\cdot(h_{LL}*F_{s}),\quad F_{s}^{lh}=h_{LH}*F_{s},$ (9)
$F_{s}^{hl}=h_{HL}*F_{s},\quad F_{s}^{hh}=h_{HH}*F_{s},$
where $\omega$ is the attenuation factor for the low-frequency components. The feature map is then reconstructed using the inverse wavelet transform:
$F_{s}^{\text{filtered}}=\text{IWT}(F_{s}^{ll},F_{s}^{lh},F_{s}^{hl},F_{s}^{hh}).$ (10)
The filtered semantic features $F_{s}^{\text{filtered}}$ are projected into the disparity space through a Conv layer with a PReLU activation function to generate the optimized disparity map. PReLUis an improved version of the ReLU activation function. It addresses the issue of ReLU’s zero output for negative inputs by introducing a learnable parameter $\alpha$ (Eq. 12). Finally, the optimized disparity map is processed using a ReLU activation function to ensure that the disparity values are non-negative:
$D_{refined}=\text{ReLU}(D+\text{PReLU}(\text{Conv}(F_{s}^{filtered}))),$ (11)
$\text{PReLU}(x)=\begin{cases}x&\text{if }x\geq 0,\\ \alpha x&\text{if }x<0.\end{cases}$ (12)
III-F Loss Function
We utilize the $smoothL_{1}$ loss(Eq. 13) to quantify the discrepancy between the predicted disparity map and the ground-truth disparity map. Specifically, we compute the loss at three critical stages of the disparity estimation pipeline: prior to cost aggregation ($d_{f}$), before disparity optimization ($d_{cg}$), and at the final output stage ($d_{dr}$).:
$L_{smooth}(x)=\begin{cases}0.5x^{2}&\text{if }|x|<1,\\ |x|-0.5&\text{otherwise}.\end{cases}$ (13)
The total loss is formulated as:
$\begin{split}\text{Loss}=w_{1}\cdot\|d_{f}-d_{gt}\|_{1}+w_{2}\cdot\|d_{cg}-d_{gt}\|_{1}+\\ w_{3}\cdot\|d_{dr}-d_{gt}\|_{1},\end{split}$ (14)
where $w_{1}$, $w_{2}$, and $w_{3}$ are the weights assigned to the losses at each respective stage(we set $w_{1}=w_{2}=w_{3}=1/3$).
Figure 3: Visualization of disparity estimation on SCARED and SERV-CT datasets. (a) and (b) are from SCARED, while (c) and (d) are from SERV-CT. When overall depth variation is minimal, as in (b), most models perform similarly. However, in scenes with significant depth transitions—indicating high-frequency components, such as between surgical instruments and tissues in (a)—LightEndoStereo provides more accurate depth estimates at boundary regions. Figures (c) and (d) show that LightEndoStereo also performs well on SERV-CT.
IV Experiments
IV-A Datasets
We trained our algorithm on the SCARED dataset and tested it on both the SCARED and SERV-CT datasets.
IV-A1 SCARED Dataset [24]
The SCARED dataset [24] is a public laparoscopy dataset from the MICCAI 2019 Endovis Challenge, captured using the da Vinci Xi surgical robot. It consists of porcine peritoneal images with a resolution of $1024\times 1280$. The dataset includes 7 training subsets and 2 test subsets. Due to significant calibration errors in datasets 4 and 5, we discarded these subsets and trained our model using the remaining training data (14,714 image pairs). The dataset 8 and 9 are testing data. We adopt the official evaluation metric: Mean Absolute Error (MAE) in mm of the depth map.
IV-A2 SERV-CT Dataset [26]
The SERV-CT dataset [26] contains 16 pairs of porcine peritoneal stereo images with a resolution of $720\times 576$. We use all 16 pairs for testing. We evaluate performance using the Mean Absolute Error (MAE) in pixels, the percentage of pixels with a disparity error greater than n pixels (Bad-n), and the percentage of pixels with an error greater than 3 pixels and greater than 5% of the ground-truth value (D1).
IV-B Implementation details
We adopted a MobileNetV4, pre-trained on ImageNet, as the decoder part branch of feature network. The model was implemented using the PyTorch framework. For optimization, we utilized the Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The learning rate for the feature network is $1\times 10^{-4}$ and $1\times 10^{-3}$ for the other parts. During training, we set the random seed to 6. For data augmentation, we applied random cropping to resize the training images to $256\times 512$. The model was trained for 100 epochs on the SCARED dataset. For evaluation, we tested the model on datasets 8 and 9 of SCARED, as well as the SERV-CT dataset. All experiments were conducted on an Ubuntu 22.04 system with 4 Nvidia RTX 2080Ti GPUs.
IV-C Evaluations on SCARED and SERV-CT
TABLE I: MAE (unit: mm) for the SCARED Test Set. Each test set comprises 5 keyframes (kf1–kf5). Lower values indicate better performance. Methods marked with (*) reference results from other papers. Bold font indicates the optimal results.
Method Dataset 8 Dataset 9
kf1 kf2 kf3 kf4 kf5 kf1 kf2 kf3 kf4 kf5 Avg.
GwcNet [16] 9.076 2.926 1.433 1.7 1.203 3.868 1.138 2.934 1.997 0.6135 2.689
PSMNet* [29] 8.96 2.77 1.43 1.83 1.42 3.99 1.08 2.82 1.95 0.56 2.68
RAFT [21] 7.921 2.362 1.721 2.167 1.88 4.363 1.087 2.872 1.983 1.549 2.79
IGEV [20] 7.955 2.393 1.697 2.265 1.938 3.893 1.174 2.895 2.504 1.25 2.796
DLNR [19] 7.848 2.529 2.654 3.209 3.418 5.374 1.913 4.324 4.82 10.25 4.634
Selective-Stereo [31] 7.537 2.409 1.668 2.31 2.264 3.86 0.958 2.838 1.775 1.018 2.664
MSDESIS-resnet32 [7] 10.14 3.467 3.306 3.889 2.099 5.428 1.636 3.378 1.768 0.6375 3.575
MSDESIS-light [7] 11.45 15.91 7.727 10.67 9.849 98.39 130.8 66.57 101.8 203.4 65.67
DCStereo* [34] 8.38 2.77 1.54 1.82 1.19 3.99 1.09 2.64 2.09 1.04 2.65
LightEndoStereo(Ours) 7.62 2.316 1.669 2.281 1.025 4.428 1.167 2.809 2.125 0.4793 2.592
TABLE II: Results of comparison experiments on the SERV-CT dataset.Methods marked with (*) reference results from other papers. Bold font indicates the optimal results.
Model MAE(pixel)$\downarrow$ D1$\downarrow$ Bad1$\downarrow$ Bad2$\downarrow$ Bad3$\downarrow$
GwcNet [16] 5.813 33.92% 79.37% 56.32% 40.93%
PSMNet* [29] 6.355 - 70.88% 49.80% 33.87%
RAFT [21] 21.500 87.50% 95.80% 91.80% 88.50%
IGEV [20] 32.895 92.91% 98.28% 96.53% 94.76%
DLNR [19] 22.030 88.97% 96.21% 92.61% 89.13%
Selective-Stereo [31] 24.103 86.36% 95.68% 91.33% 87.01%
MSDESIS-resnet32 [7] 5.742 28.26% 78.81% 55.44% 37.79%
MSDESIS-light [7] 7.065 41.35% 81.09% 63.17% 49.22%
LightEndoStereo(Ours) 2.367 14.34% 58.59% 31.52% 18.49%
We evaluated our algorithm against several methods for natural image stereo matching and two algorithms for endoscopic data. Specifically, GwcNet [16], PSMNet [29], RAFT [35], IGEV [20], DLNR [19], and Selective-Stereo [31] are methods designed for natural images, with DLNR and Selective-Stereo focusing on optimizing high-frequency details in disparity maps. For these natural image methods, except for PSMNet, we retrained them on the SCARED dataset. For PSMNet, we directly cited the results from [34, 3]. MSDESIS [7] and DCStereo [34] are algorithms tailored for endoscopic images. DCStereo’s evaluation methodology on SCARED aligns with ours, so we directly cited their results. However, since DCStereo was not tested on the SERV-CT dataset and its code is not publicly available, we omitted their results on SERV-CT. MSDESIS was trained on the SCARED dataset, and we directly loaded their pre-trained weights. Given that our maximum disparity is set to 192, whereas MSDESIS uses a maximum disparity of 320, the results we reproduced differ slightly from those reported in their paper. The evaluation results on SCARED are shown in Table I, and the results on SERV-CT are shown in Table II. LightEndoStereo achieved SOTA performance on both datasets and demonstrated excellent generalization on SERV-CT.
IV-D Runtime Evaluation
TABLE III: Quantitative results of model computational performance. Note that only methods with an average MAE < 3mm on the SCARED dataset are included.Methods marked with (*) reference results from other papers. Bold font indicates the optimal results.
Model Params(M) FLOPs(T) Runtimes(ms)
GwcNet [16] 6.43 2.466 26.4
PSMNet* [29] 3.672 - 225
RAFT [21] 11.11 3.757 468.6
IGEV [20] 12.5 3.288 434.5
Selective-Stereo [31] 13.141 4.062 516.52
DCStereo* [34] 3.404 - 191
LightEndoStereo(Ours) 11.094 0.846 23.38
To evaluate the runtime performance of our model, we conducted tests using a single Nvidia RTX 2080Ti GPU, paired with an Intel(R) Xeon(R) Gold 6130T CPU @ 2.10GHz. We sampled 100 images ($1280\times 1024$) from dataset 8 of the SCARED dataset for testing and used the average inference time as the evaluation metric. The results are shown in Table III. Although the MSDESIS-Light model has the smallest footprint, its matching accuracy is poor. In contrast, LightEndoStereo achieved a real-time matching speed of 23.38 ms per frame.
Figure 4: Visualization of disparity optimization in high-frequency regions. The top row shows the depth maps, while the bottom row displays the corresponding error maps.
TABLE IV: Ablation Study Results. We use the average MAE on the SCARED dataset as the accuracy evaluation metric. When ablating the 3D UNet-MCA, we employ the widely-used 3D Hourglass aggregation network.
3D UNet-MCA HFDO MAE(mm) Runtime(ms) Params(M) FLOPs(T)
2.718 25.412ms 13.0479 1.7552
$\checkmark$ 2.645 25.972ms 13.0773 1.7579
2.621 22.655ms 11.0649 0.8435
$\checkmark$ 2.592 23.386ms 11.0943 0.8461
IV-E Ablation Study
We conducted detailed ablation studies on the proposed 3D MCA and HFDO modules (see Table IV). When ablating the 3D UNet-MCA aggregation network, we employ the widely-used 3D Hourglass aggregation network [16]. The 3D Hourglass aggregation network is a stacked symmetric multi-scale network that features a large number of parameters and high computational complexity. As shown in Table III, the 3D UNet-MCA not only reduces the number of parameters and computational complexity but also enhances disparity estimation accuracy. As Figure 4 illustrates, the HFDO module significantly reduces matching errors at tissue boundaries, leading to notable improvements in both the Stacked 3D Hourglass and 3D UNet-MCA configurations.
V Conclusions and limitations
We introduce LightEndoStereo, a novel lightweight real-time stereo matching algorithm tailored for endoscopic images. Our approach integrates a coordinate attention mechanism and the Mamba block to achieve efficient cost aggregation via a lightweight 3D UNet. Moreover, by enhancing high-frequency features in the wavelet domain, we optimize disparity estimation, effectively mitigating large matching errors at tissue boundaries—a common challenge in endoscopic stereo matching. During validation, LightEndoStereo exhibited strong generalization on the SERV-CT dataset, a potential that warrants further exploration. Acknowledging the significance of generalizable stereo matching algorithms, we intend to delve deeper into this area in our future research.

### v3

@@
-LightEndoStereo: A Real-time Lightweight Stereo Matching Method for Endoscopy Images
+Robust Real-Time Endoscopic Stereo Matching under Fuzzy Tissue Boundaries
 Abstract
-Real-time acquisition of accurate depth of scene is essential for automated robotic minimally invasive surgery, and stereo matching with binocular endoscopy can generate such depth.
-However, existing algorithms struggle with ambiguous tissue boundaries and real-time performance in prevalent high-resolution endoscopic scenes.
-We propose LightEndoStereo, a lightweight real-time stereo matching method for endoscopic images.
-We introduce a 3D Mamba Coordinate Attention module to streamline the cost aggregation process by generating position-sensitive attention maps and capturing long-range dependencies across spatial dimensions using the Mamba block.
-Additionally, we introduce a High-Frequency Disparity Optimization module to refine disparity estimates at tissue boundaries by enhancing high-frequency information in the wavelet domain.
-Our method is evaluated on the SCARED and SERV-CT datasets, achieving state-of-the-art matching accuracy and a real-time inference speed of 42 FPS.
-The code is available at https://github.com/Sonne-Ding/LightEndoStereo.
-I Introduction
+Real-time acquisition of accurate scene depth is essential for automated robotic minimally invasive surgery.
+Stereo matching with binocular endoscopy can provide this depth information.
+However, existing stereo matching methods, designed primarily for natural images, often struggle with endoscopic images due to fuzzy tissue boundaries and typically fail to meet real-time requirements for high-resolution endoscopic image inputs.
+To address these challenges, we propose RRESM, a real-time stereo matching method tailored for endoscopic images.
+Our approach integrates a 3D Mamba Coordinate Attention module that enhances cost aggregation through position-sensitive attention maps and long-range spatial dependency modeling via the Mamba block, generating a robust cost volume without substantial computational overhead.
+Additionally, we introduce a High-Frequency Disparity Optimization module that refines disparity predictions near tissue boundaries by amplifying high-frequency details in the wavelet domain.
+Evaluations on the SCARED and SERV-CT datasets demonstrate state-of-the-art matching accuracy with a real-time inference speed of 42 FPS.
+The code is available at https://github.com/Sonne-Ding/RRESM.
+1 Introduction
 Minimally invasive surgery (MIS) has become a preferred surgical approach due to its reduced invasiveness and faster recovery times [21].
 Endoscopy can provide essential visual guidance in MIS.
-However, endoscopy faces inherent limitations, including a restricted field of view, lack of tactile feedback, and diminished spatial awareness [2, 3, 4].
-To address these challenges, computer-assisted intervention (CAI) techniques have been developed to extract spatial depth information from endoscopic images, with depth estimation emerging as a key focus [3].
+Although endoscopy provides essential visual guidance in MIS, it faces inherent limitations—such as a restricted field of view, lack of tactile feedback, and diminished spatial awareness [4, 27].
+To overcome these challenges, computer-assisted intervention techniques have been developed to extract spatial depth information from endoscopic images, with depth estimation emerging as a key focus [27].
 Endoscopic stereo matching is the method to obtain tissue depth information from binocular images [34, 2].
@@
 Although there have been numerous studies on stereo matching in the domain of natural images, endoscopic stereo matching still faces some challenges due to its unique application scenarios.
-Tissue boundaries in endoscopic images are blurred due to smooth transitions [7], making accurate depth estimation difficult.
-Meanwhile, the increasing adoption of high-definition endoscopes adds challenges for stereo matching to maintain computational efficiency and real-time performance.
+Tissue boundaries in endoscopic images are blurred by low-contrast, gradual transitions [20], making accurate depth estimation difficult.
+Meanwhile, the increasing adoption of high-definition endoscopes places additional burden on stereo matching to preserve computational efficiency and real-time performance.
 Therefore, developing an efficient, accurate, and robust endoscopic stereo matching method is essential.
-Figure 1: Overview of LightEndoStereo.
-The figure presents the average MAE and single-frame inference speed of various methods on the SCARED dataset.
-Our method achieves state-of-the-art (SOTA) matching accuracy while maintaining high inference speed.
-Since Jure Zbontar [8] first introduced convolutional neural networks (CNNs) to stereo matching, numerous deep learning-based methods have been proposed in this field.
-Compared to traditional methods, deep learning approaches, leveraging complex network architectures, adaptive feature extraction, and refined optimization capabilities, have achieved significantly higher matching accuracy [9, 10, 11, 12, 13].Currently, there are two mainstream approaches in stereo matching: those based on 3D convolutions [5, 8, 14, 15, 16, 9, 17, 18] and those based on iterative optimization [19, 20, 21, 22].
-Methods based on 3D convolutions can effectively encode essential geometric information for matching.
-However, they also introduce increased computational complexity [23].
-On the other hand, iterative optimization frameworks employ recurrent neural networks (RNNs) to iteratively refine disparity estimation [20], balancing matching accuracy and computational performance by adjusting the number of iterations.When processing high-resolution endoscopic images (e.g., $1024\times 1280$ in the SCARED dataset [24]), these iterative methods often require multiple iterations to achieve satisfactory results.
-Due to the lack of GPU acceleration, their inference speed lags behind convolution-based methods and fails to meet real-time requirements.
-[…]
-Our main contributions are summarized as follows:
-  • We explore stereo matching methods for real-time depth information acquisition via binocular endoscopy in automated minimally invasive surgery, and propose LightEndoStereo, a lightweight real-time stereo matching algorithm that enables accurate disparity estimation, particularly in high-frequency regions such as tissue boundaries and surgical instrument edges.
-  • To achieve high-quality aggregation of matching costs in 3D space with minimal computational overhead, we propose the 3D Mamba Coordinate Attention (MCA) module.
-This module incorporates spatial information into channel attention and efficiently captures long-range correlations across the entire 3D space.
-  • To minimize disparity estimation errors at boundaries, which are particularly evident in endoscopic images, we propose a High-Frequency Disparity Optimization (HFDO) module.
-This module enhances high-frequency components of context features in the wavelet domain to improve disparity estimation at tissue and instrument boundaries.
-II RELATED WORK
-II-A Real-time Stereo Matching
-For of real-time stereo matching, there are primarily two categories: CNN-based cost aggregation methods and iterative optimization-based methods The use of CNNs in stereo matching traces back to the nascent stages of deep learning.
+Figure 1: Depth error maps near fuzzy boundaries on the SCARED dataset.
+Brighter regions indicate larger depth errors.
+Our method performs well in estimating depth around fuzzy tissue boundaries, outperforming the state-of-the-art natural image methods (e.g., GwcNet and IGEV).
+Since Zbontar et al.
+[31] first introduced convolutional neural networks (CNNs) into stereo matching, numerous deep learning-based methods have been proposed in this field.
+Compared to traditional approaches, deep models—leveraging complex architectures, adaptive feature extraction, and learned optimization—have achieved significantly higher matching accuracy [6, 33, 12, 23].
+Currently, stereo matching methods primarily fall into two categories: those based on 3D convolutions [34, 31, 24, 10, 32, 29], and those based on iterative optimization [35, 28, 16, 15].
+3D convolution-based approaches effectively encode geometric information but incur substantial computational overhead [30].
+Iterative optimization methods, typically employing recurrent neural networks (RNNs) to refine disparity estimations over multiple steps [28], offer a trade-off between accuracy and efficiency by adjusting the number of iterations.
+However, they still exhibit high latency when processing high-resolution endoscopic images.
+Moreover, most of these models are designed for natural image domains and generalize poorly to endoscopic scenes.
+As shown in Fig.
+1, both the 3D CNN-based GwcNet [10] and the iterative method IGEV [28] produce large depth errors near fuzzy tissue boundaries.
+To enable real-time stereo matching for robotic applications in endoscopic scenes with fuzzy tissue boundaries, we propose RRESM, a lightweight and efficient stereo matching framework.
+RRESM employs MobileNetV4 [22] as a compact feature extractor to reduce computational overhead while preserving representational capacity.
+For cost aggregation, we introduce a novel attention-based module, the 3D Mamba Coordinate Attention (MCA) Module.
+Instead of computing attention over the entire 3D cost volume, MCA performs axis-wise attention in 1D, significantly reducing computational complexity.
+It further leverages the Mamba block [5] to model long-range dependencies along all three spatial dimensions.
+This design enables effective cost aggregation with minimal computational overhead.
+Experimental results show that MCA significantly outperforms traditional architectures such as the Stacked Hourglass Network [20], which relies heavily on stacked 3D convolutions.
+To further refine disparity maps, especially around tissue boundaries, we propose the High-Frequency Disparity Optimization (HFDO) module.
+This module uses the Haar wavelet transform to decompose contextual features into low- and high-frequency components.
+It enhances high-frequency signals and reconstructs features via inverse wavelet transform, thereby enriching the disparity representation with fine structural details.
+The refined features are then projected into disparity space to enhance depth estimation in high-frequency regions.
+Our method demonstrates superior performance in depth estimation near fuzzy tissue boundaries, as shown in Fig.1.
+We evaluate RRESM on the SCARED [1] and SERV-CT [7] datasets, demonstrating both accuracy and real-time performance.
+Our method achieves the best average MAE of $2.592\,\text{mm}$ on the SCARED dataset, with an inference time of $23.38\,\text{ms}$ per frame at a resolution of $1024\times 1280$.
+2 Related Works
+2.1 Fuzzy Boundaries Optimization in Stereo Matching
+Boundary regions in images are typically characterized by high-frequency components, which pose challenges for accurate stereo matching.
+Several recent works have focused on improving disparity estimation in such regions by enhancing high-frequency detail prediction [35, 26, 18].
+DLNR [35] points out that the tight coupling between the update matrix and hidden state transition in the GRU module used by RAFT can lead to degraded performance in high-frequency regions.
+To mitigate this, DLNR replaces GRUs with LSTMs (Long Short-Term Memory), achieving improved accuracy in disparity prediction near edges.Selective-Stereo [26] further observes that GRUs with fixed receptive fields are limited in their ability to capture both high-frequency edge information and low-frequency texture information.
+To address this, it introduces the Selective Recurrent Unit (SRU), which incorporates multi-scale receptive fields into the GRU framework.
+This design allows the network to adaptively process features across multiple frequency bands, enhancing disparity estimation performance in both boundary and low-frequency texture regions.
+2.2 Real-time Stereo Matching
+For real-time stereo matching, there are primarily two categories: CNN-based cost aggregation methods and iterative optimization-based methods The use of CNNs in stereo matching traces back to the nascent stages of deep learning.
 These methods predominantly utilize CNNs to extract features and construct a cost volume for disparity estimation.
@@
 Similarly, DLNR [35] employs an LSTM (Long-Short-Term Memory) structure to specifically optimize disparity estimates in high-frequency regions, thereby further bolstering the overall performance of the models.
-II-B High-Frequency Optimization in Stereo Matching
-Several works have focused on optimizing stereo matching for high-frequency components [19, 31, 32].The DLNR [19] highlights that the coupling between the update matrix and the hidden state transfer in the GRU module used by RAFT can lead to inaccuracies in high-frequency detail prediction.
-To address this, DLNR introduced an LSTM (Long-Short-Term Memory) structure to replace the GRU, thereby achieving better high-frequency disparity prediction performance.
-The Selective-Stereo [31] highlighted that GRUs with fixed receptive fields struggle to capture both high-frequency information in edges and low-frequency information in smooth regions simultaneously.
-To overcome this limitation, Selective-Stereo proposed the Selective Recurrent Unit (SRU), which incorporates multi-scale receptive fields into the GRU design.
-This approach enables the model to optimize across multiple frequency bands, improving the overall accuracy of disparity prediction.
-Figure 2: (a) Architecture of LightEndoStereo.
-The Feature Net is a UNet-like structure with an encoder branch based on MobileNetV4.
-The group-wise correlation cost volume is utilized.
-(b) Mamba Coordinate Attention module.
-Attention is extracted separately across the $H$, $W$, and $D$ dimensions, concatenated along the channel dimension, and then passed through a Bi-Mamba2 layer.
-Bi-Mamba2 layer extracts dependencies between features in both forward and backward directions across all spatial dimensions ($H+W+D$).
-The resulting position-sensitive attention maps are used to re-weight the feature map.
-(c) Wavelet Transform Refine module.
-The context feature map is decomposed into low-frequency (LL) and high-frequency (LH, HL, HH) components.
-The LL components are attenuated by a parameter $w$.
-III METHOD
-In this section, we present the methodology of LightEndoStereo, a lightweight stereo matching algorithm designed for high-precision disparity estimation while maintaining efficient computational performance.
-We integrate a lightweight feature extraction network, an MCA module for optimized cost aggregation, and an HFDO module to refine disparities in high-frequency regions.
-The overall architecture is shown in Fig.
+3 Method
+The overall architecture of RRESM is illustrated in Fig.
 2.
-III-A Feature Extraction
-Multi-scale information is essential for stereo matching.
-Shallow features capture rich texture and geometric details, while deep features provide important abstract semantic information.
-As so, we propose a lightweight multi-scale feature extraction network based on the U-Net architecture.
-The encoder branch utilizes MobileNetV4, an efficient lightweight neural network optimized for mobile devices.
-In the decoder stage, we employ 2D convolutions and skip connections between adjacent scales to progressively restore multi-scale information.
-Given a stereo pair $I_{l},I_{r}\in\mathbb{R}^{3\times H\times W}$, we utilize the MobileNetV4 structure [25] as the downsampling branch of the U-Net multi-scale feature extraction framework.
-In the upsampling branch, features at scales of $1/16$, $1/8$, and $1/4$ are fused, resulting in the final output features at a $1/4$ scale of the original image, $F_{l},F_{r}\in\mathbb{R}^{C\times H/4\times W/4}$.
-The feature $F_{l}$ extracted from the encoder branch serves as the context feature for subsequent disparity optimization.
-III-B Cost Volume Computation
+The feature extraction backbone is based on MobileNetV4, providing a lightweight and efficient representation.
+The proposed MCA module is incorporated into a simplified 3D U-Net to perform cost aggregation in the feature space.
+Subsequently, the HFDO module is applied in the frequency domain to refine disparity predictions, particularly near boundary regions.
+Detailed descriptions of each component are provided in the following sections.
+Figure 2: Overall architecture of RRESM.
+The Feature Net adopts a U-Net-like structure, with a frozen encoder based on MobileNetV4 and a trainable decoder.
+A correlation cost volume is constructed using group-wise correlation.
+The MCA module is embedded within a simplified 3D CNN to enhance cost aggregation.
+The HFDO module takes deep features from the encoder as contextual information, applies a wavelet-based filter, and enhances high-frequency components in the disparity map.
+Details of the MCA and HFDO modules are shown in Fig.3 4, respectively.
+3.1 Feature Extraction
+Multi-scale feature information is crucial for stereo matching: shallow features capture rich textures and fine geometric details, while deeper features capture higher-level semantics.
+To leverage both, we adopt a U-Net-like structure, where the downsampling branch is based on a frozen MobileNetV4, and the upsampling branch is trainable and gradually integrates multi-scale features to recover spatial resolution.
+Formally, given a stereo pair $I_{l},I_{r}\in\mathbb{R}^{3\times H\times W}$, we use the MobileNetV4 backbone [22] as the downsampling path in our multi-scale feature extraction framework.
+In the upsampling path, features from resolutions of $1/16$, $1/8$, and $1/4$ are progressively fused, producing final feature maps at $1/4$ resolution: $F_{l},F_{r}\in\mathbb{R}^{C\times H/4\times W/4}$.
+The feature $F_{l}$ extracted from the downsampling branch is also used as context input for the subsequent disparity refinement module.
+3.2 Cost Volume Computation
 We employ a Group-wise Correlation cost volume [10] to compute the matching cost between the extracted paired features.
@@
 $CV(g,d,x,y)=\frac{1}{C/g_{n}}\langle F_{l}^{g}(x,y),F_{r}^{g}(x-d,y)\rangle,$ (1)
-[…]
-III-C 3D Mamba Coordinate Attention Guided Cost Aggregation
-We propose a novel 3D Mamba Coordinate Attention module.
-This module extends the coordinate attention framework by embedding positional information into channel attention.
-Moreover, to capture long-range interactions among channel features across the disparity, width, and height dimensions, we introduce a bidirectional Mamba2 block.
-Specifically, for the cost volume $CV\in\mathbb{R}^{g_{n}\times D\times H\times W}$, we treat the disparity dimension $D$ as an additional spatial dimension.
-We employ three pooling kernels $(H,W)$, $(D,H)$, and $(D,W)$ to perform channel encoding along the horizontal, vertical, and disparity directions, respectively.
-For a given 3D feature $\mathcal{F}\in\mathbb{R}^{C\times X\times Y\times Z}$ (where $Z$ dimension corresponds to the disparity $D$ in this work), the channel-wise attention extraction for the $c$-th channel along the three directions is formulated as:
-$\mathcal{Z}^{x}_{c}(x)$ $=\frac{1}{Y\cdot Z}\max_{y,z}\mathcal{F}(c,x,y,z),$ (2)
-$\mathcal{Z}^{y}_{c}(y)$ $=\frac{1}{X\cdot Z}\max_{x,z}\mathcal{F}(c,x,y,z),$
-$\mathcal{Z}^{z}_{c}(z)$ $=\frac{1}{X\cdot Y}\max_{x,y}\mathcal{F}(c,x,y,z).$
-Subsequently, we introduce non-linearity via a sigmoid function and employ the Mamba2 module to capture long-range dependencies across all spatial dimensions.
-The refined spatial attention is computed as:
-[…]
-$A_{x},A_{y},A_{z}=\text{split}(A_{\text{refined}}),$
-where $\delta(\cdot)$ denotes the sigmoid function.
-Finally, the refined attention maps $A_{x}$, $A_{y}$, and $A_{z}$ are applied to the original feature map $\mathcal{F}$ to obtain the optimized feature map:
+where $CV$ denotes the cost volume, $F_{*}^{g}$ represents the feature vectors within group $g$ ($g\in[0,g_{n}-1]$), and $\langle\cdot,\cdot\rangle$ denotes the inner product.
+3.3 3D Mamba Coordinate Attention Guided Cost Aggregation
+Figure 3: (a) Architecture of the MCA module.
+Attention is computed independently along the $H$, $W$, and $D$ dimensions, concatenated along the channel axis, and passed through a Bi-Mamba2 layer.
+(b) Implementation of the Bi-Mamba2 layer in MCA.
+A bidirectional scan is performed over the concatenated axis-pooled features in both forward and backward directions.
+(c) Re-weighting operation.
+The resulting position-sensitive attention maps assign a unique weight to each coordinate in the cost volume.
+We propose the 3D Mamba Coordinate Attention (MCA) module to perform efficient, position-sensitive cost aggregation, as illustrated in Fig.
+3.
+MCA enhances the conventional coordinate attention mechanism by embedding axis-specific positional information and explicitly modeling long-range dependencies across all three spatial dimensions—disparity (D), height (H), and width (W).
+Recently, Mamba has emerged as a state-space model (SSM)-based alternative to transformers, offering strong sequence modeling capabilities with linear time and space complexity [5].
+Unlike attention mechanisms that compute pairwise token interactions with quadratic cost, Mamba leverages structured state-space dynamics to capture global dependencies via fast kernel-based operations.
+This makes it well suited for real-time or high-resolution tasks, where transformers typically incur prohibitive computational overhead.
+Moreover, Mamba excels at capturing long-range spatial interactions, which aligns well with the requirements of 3D cost aggregation in stereo matching.
+To incorporate spatial priors efficiently, we extend Mamba into a 3D axis-aware attention mechanism.
+Specifically, we treat the disparity dimension $D$ as an additional spatial axis.
+Given a cost volume $\mathcal{F}\in\mathbb{R}^{C\times D\times H\times W}$, we extract axis-aligned attention descriptors by applying global pooling over orthogonal spatial planes:
+$\mathcal{Z}^{x}_{c}(x)$ $=\frac{1}{H\cdot D}\max_{y,z}\mathcal{F}(c,x,y,z),$ (2)
+$\mathcal{Z}^{y}_{c}(y)$ $=\frac{1}{W\cdot D}\max_{x,z}\mathcal{F}(c,x,y,z),$
+$\mathcal{Z}^{z}_{c}(z)$ $=\frac{1}{W\cdot H}\max_{x,y}\mathcal{F}(c,x,y,z),$
+where $\mathcal{Z}^{x}$, $\mathcal{Z}^{y}$, and $\mathcal{Z}^{z}$ denote the axis-specific pooled descriptors. These vectors are passed through a sigmoid activation and concatenated as:
+$A_{\text{cat}}=\text{concat}\left[\delta(\mathcal{Z}^{x}),\delta(\mathcal{Z}^{y}),\delta(\mathcal{Z}^{z})\right],$ (3)
+where $\delta(\cdot)$ is the sigmoid function.
+The original Mamba (Mamba1) [9] performs unidirectional sequence modeling along a given axis, which may lead to directional bias.
+To enhance symmetry and spatial coherence, we employ the Mamba2, which applies bidirectional sequential modeling along each axis.
+This enables the model to aggregate contextual cues both forward and backward across disparity, height, and width dimensions, which is particularly beneficial near fuzzy anatomical boundaries.
+The refined attention maps are obtained as:
+$A_{\text{refined}}=\text{Mamba2}(A_{\text{cat}}),\quad A_{x},A_{y},A_{z}=\text{split}(A_{\text{refined}}),$ (4)
+and applied to the original cost volume by broadcasted element-wise multiplication:
 $\mathcal{F}_{\text{refined}}=\mathcal{F}\odot A_{x}\odot A_{y}\odot A_{z},$ (5)
-where $\odot$ denotes broadcast $A_{*}$ and then performs element-wise multiplication.
-The proposed 3D MCA module is embedded into a light 3 scale 3D UNet cost aggregation network to enhance the model’s ability to capture long-range dependencies across the entire spatial domain.
-We refer to this aggregation network as 3D UNet-MCA in the following text.
-III-D Disparity Computation
-To achieve differentiable disparity computation and generate sub-pixel disparity estimates with higher precision, we utilize the soft argmax operation to calculate the disparity [28].
-Specifically, the matching cost volume $CV$ is first processed by reducing its feature dimension to 1.
-Subsequently, the spatial scale is restored through bilinear interpolation, resulting in the final cost volume $CV\in\mathbb{R}^{D\times H\times W}$.
-The operations are formulated as follows:
-$CV=\text{Interpolate}(\text{Conv}(CV)),$ (5)
-where Conv denotes the convolution operation that compresses the feature dimension, and Interpolate represents the bilinear interpolation for spatial scale recovery.
-The disparity at each pixel location $(h,w)$ is then computed using the soft argmax operation:
-$d(h,w)=\sum_{i=0}^{D-1}i\times\sigma(CV(i,h,w)),$ (6)
-where $\sigma$ is the softmax function.
-III-E High Frequency Disparity Optimization
-In endoscopic images, depth variations are relatively gradual in most regions.
-This characteristic poses challenges for the model in predicting depth in high-frequency regions, while the boundaries and blood vessels are not prominent in the images.
-To enhance the optimization of high-frequency regions in the disparity optimization process, while avoiding introducing noise to the well-predicted low-frequency regions, we employ a wavelet module.
-This module decomposes the context feature $F_{s}$ obtained from $F_{l}$ into low-frequency components(LL) and high-frequency components in horizontal(LH), vertical(HL), and diagonal directions(HH).
-The low-frequency components are then attenuated by a scale $\omega$, and the inverse wavelet transform is applied to obtain the refined context feature.
-The refined context feature specifically targets the optimization of high-frequency detail regions in the disparity map.
+where $\odot$ denotes Hadamard (element-wise) product.
+The refined feature volume $\mathcal{F}_{\text{refined}}$ is then passed to a lightweight 3D U-Net cost aggregation network, improving global spatial reasoning while maintaining low computational complexity.
+3.4 High Frequency Disparity Optimization
+Figure 4: Wavelet Transform Refine module.
+The context feature map is decomposed into low-frequency (LL) and high-frequency (LH, HL, $H$) components.
+The $L$ components are attenuated by a parameter $w$.
+To further improve disparity estimation near fuzzy boundaries, we introduce a wavelet-based module, HFDO, which refines the disparity map in the frequency domain.
+This module decomposes the context feature $F_{s}$, extracted from $F_{l}$, into low-frequency components (LL) and high-frequency components in the horizontal (LH), vertical (HL), and diagonal (HH) directions via a Haar wavelet transform.
+To emphasize high-frequency structures, the low-frequency component is attenuated by a scaling factor $\omega$, and the inverse wavelet transform is subsequently applied to reconstruct a refined context feature.
+This refined representation specifically enhances high-frequency details, improving disparity prediction in regions with ambiguous or fuzzy tissue boundaries.
 The context features are derived as follows:
@@
 where Conv denotes a 2D $3\times 3$ convolution. The 2D Haar wavelet convolution kernels used are [8]:
-$h_{LL}=\frac{1}{2}\begin{bmatrix}1&1\\ 1&1\end{bmatrix},\quad h_{LH}=\frac{1}{2}\begin{bmatrix}1&-1\\ 1&-1\end{bmatrix},$ (8)
-$h_{HL}=\frac{1}{2}\begin{bmatrix}1&1\\ -1&-1\end{bmatrix},h_{HH}=\frac{1}{2}\begin{bmatrix}1&-1\\ -1&1\end{bmatrix}.$
+$h_{LL}=\frac{1}{2}\begin{bmatrix}1&1\\ 1&1\end{bmatrix},\ h_{LH}=\frac{1}{2}\begin{bmatrix}1&-1\\ 1&-1\end{bmatrix},\ h_{HL}=\frac{1}{2}\begin{bmatrix}1&1\\ -1&-1\end{bmatrix},\ h_{HH}=\frac{1}{2}\begin{bmatrix}1&-1\\ -1&1\end{bmatrix}$.
+(7)
 The decomposition of the feature map $F_{s}$ is performed as follows:
-$F_{s}^{ll}=\omega\cdot(h_{LL}*F_{s}),\quad F_{s}^{lh}=h_{LH}*F_{s},$ (9)
-$F_{s}^{hl}=h_{HL}*F_{s},\quad F_{s}^{hh}=h_{HH}*F_{s},$
+$F_{s}^{ll}=\omega\cdot(h_{LL}*F_{s}),\ F_{s}^{lh}=h_{LH}*F_{s},\ F_{s}^{hl}=h_{HL}*F_{s},\ F_{s}^{hh}=h_{HH}*F_{s},$ (8)
 where $\omega$ is the attenuation factor for the low-frequency components. The feature map is then reconstructed using the inverse wavelet transform:
 $F_{s}^{\text{filtered}}=\text{IWT}(F_{s}^{ll},F_{s}^{lh},F_{s}^{hl},F_{s}^{hh}).$ (9)
-The filtered semantic features $F_{s}^{\text{filtered}}$ are projected into the disparity space through a Conv layer with a PReLU activation function to generate the optimized disparity map.
-PReLUis an improved version of the ReLU activation function.
-It addresses the issue of ReLU’s zero output for negative inputs by introducing a learnable parameter $\alpha$ (Eq.
-12).
+The filtered semantic features $F_{s}^{\text{filtered}}$ are projected into the disparity space through a Conv layer with a PReLU [11] activation function to generate the optimized disparity map.
 Finally, the optimized disparity map is processed using a ReLU activation function to ensure that the disparity values are non-negative:
 $D_{refined}=\text{ReLU}(D+\text{PReLU}(\text{Conv}(F_{s}^{filtered}))),$ (10)
-$\text{PReLU}(x)=\begin{cases}x&\text{if }x\geq 0,\\ \alpha x&\text{if }x<0.\end{cases}$ (12)
-III-F Loss Function
+3.5 Loss Function
 We utilize the $smooth\ L_{1}$ loss(Eq.
-13) to quantify the discrepancy between the predicted disparity map and the ground-truth disparity map.
+11) to quantify the discrepancy between the predicted disparity map and the ground-truth disparity map.
 Specifically, we compute the loss at three critical stages of the disparity estimation pipeline: prior to cost aggregation ($d_{f}$), before disparity optimization ($d_{cg}$), and at the final output stage ($d_{dr}$).:
@@
 The total loss is formulated as:
-$\begin{split}\text{Loss}=w_{1}\cdot\|d_{f}-d_{gt}\|_{1}+w_{2}\cdot\|d_{cg}-d_{gt}\|_{1}+\\ w_{3}\cdot\|d_{dr}-d_{gt}\|_{1},\end{split}$ (14)
+$\begin{split}\text{Loss}=w_{1}\cdot\|d_{f}-d_{gt}\|_{1}+w_{2}\cdot\|d_{cg}-d_{gt}\|_{1}+w_{3}\cdot\|d_{dr}-d_{gt}\|_{1},\end{split}$ (12)
 where $w_{1}$, $w_{2}$, and $w_{3}$ are the weights assigned to the losses at each respective stage(we set $w_{1}=w_{2}=w_{3}=1/3$).
-Figure 3: Visualization of disparity estimation on SCARED and SERV-CT datasets.
-(a) and (b) are from SCARED, while (c) and (d) are from SERV-CT.
-When overall depth variation is minimal, as in (b), most models perform similarly.
-However, in scenes with significant depth transitions—indicating high-frequency components, such as between surgical instruments and tissues in (a)—LightEndoStereo provides more accurate depth estimates at boundary regions.
-Figures (c) and (d) show that LightEndoStereo also performs well on SERV-CT.
-IV Experiments
-IV-A Datasets
-We trained our algorithm on the SCARED dataset and tested it on both the SCARED and SERV-CT datasets.
-IV-A1 SCARED Dataset [24]
+4 Experiments
+4.1 Datasets
+We evaluate our method on two publicly available dataset:
 The SCARED dataset [1] is a public laparoscopy dataset from the MICCAI 2019 Endovis Challenge, captured using the da Vinci Xi surgical robot.
@@
 We adopt the official evaluation metric: Mean Absolute Error (MAE) in mm of the depth map.
-IV-A2 SERV-CT Dataset [26]
 The SERV-CT dataset [7] contains 16 pairs of porcine peritoneal stereo images with a resolution of $720\times 576$.
@@
 We evaluate performance using the Mean Absolute Error (MAE) in pixels, the percentage of pixels with a disparity error greater than n pixels (Bad-n), and the percentage of pixels with an error greater than 3 pixels and greater than 5% of the ground-truth value (D1).
-IV-B Implementation details
-We adopted a MobileNetV4, pre-trained on ImageNet, as the decoder part branch of feature network.
-The model was implemented using the PyTorch framework.
-For optimization, we utilized the Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$.
-The learning rate for the feature network is $1\times 10^{-4}$ and $1\times 10^{-3}$ for the other parts.
-During training, we set the random seed to 6.
+4.2 Implementation Details
+Our implementation is based on the PyTorch framework.
+We utilized the Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$.
+The learning rate is set to $1\times 10^{-4}$.
 For data augmentation, we applied random cropping to resize the training images to $256\times 512$.
 The model was trained for 100 epochs on the SCARED dataset.
-For evaluation, we tested the model on datasets 8 and 9 of SCARED, as well as the SERV-CT dataset.
+For evaluation, we tested the model on datasets 8 and 9 of the SCARED dataset, as well as the entire SERV-CT dataset.
 All experiments were conducted on an Ubuntu 22.04 system with 4 Nvidia RTX 2080Ti GPUs.
-IV-C Evaluations on SCARED and SERV-CT
-TABLE I: MAE (unit: mm) for the SCARED Test Set.
+4.3 Evaluations on SCARED and SERV-CT
+We evaluated our algorithm against a set of state-of-the-art stereo matching methods, including both those designed for natural images and those specifically developed for endoscopic data.
+In particular, GwcNet [10], PSMNet [3], RAFT [17], IGEV [28], DLNR [35], and Selective-Stereo [26] are representative natural image methods.
+Among them, DLNR and Selective-Stereo explicitly focus on optimizing high-frequency details in disparity estimation.
+Except for PSMNet, which we report using results cited from [13, 27], all other models were retrained on the SCARED dataset under a consistent training pipeline.
+Table 1: Results of comparison experiments on the SERV-CT dataset.
+[…]
+GwcNet 5.813 33.92% 79.37% 56.32% 40.93%
+PSMNet* 6.355 - 70.88% 49.80% 33.87%
+RAFT 21.500 87.50% 95.80% 91.80% 88.50%
+IGEV 32.895 92.91% 98.28% 96.53% 94.76%
+DLNR 22.030 88.97% 96.21% 92.61% 89.13%
+Selective-Stereo 24.103 86.36% 95.68% 91.33% 87.01%
+MSDESIS-full 5.742 28.26% 78.81% 55.44% 37.79%
+MSDESIS-light 7.065 41.35% 81.09% 63.17% 49.22%
+RRESM(ours) 2.367 14.34% 58.59% 31.52% 18.49%
+Table 2: MAE (unit: $m$) for the SCARED Test Set.
 Each test set comprises 5 keyframes (kf1–kf5).
@@
 Methods marked with (*) reference results from other papers.
-Bold font indicates the optimal results.
+Bold indicates the best performance, and underlined values indicate the second-best.
 Method Dataset 8 Dataset 9
 kf1 kf2 kf3 kf4 kf5 kf1 kf2 kf3 kf4 kf5 Avg.
-GwcNet [16] 9.076 2.926 1.433 1.7 1.203 3.868 1.138 2.934 1.997 0.6135 2.689
-PSMNet* [29] 8.96 2.77 1.43 1.83 1.42 3.99 1.08 2.82 1.95 0.56 2.68
-RAFT [21] 7.921 2.362 1.721 2.167 1.88 4.363 1.087 2.872 1.983 1.549 2.79
-IGEV [20] 7.955 2.393 1.697 2.265 1.938 3.893 1.174 2.895 2.504 1.25 2.796
-DLNR [19] 7.848 2.529 2.654 3.209 3.418 5.374 1.913 4.324 4.82 10.25 4.634
-Selective-Stereo [31] 7.537 2.409 1.668 2.31 2.264 3.86 0.958 2.838 1.775 1.018 2.664
-MSDESIS-resnet32 [7] 10.14 3.467 3.306 3.889 2.099 5.428 1.636 3.378 1.768 0.6375 3.575
-MSDESIS-light [7] 11.45 15.91 7.727 10.67 9.849 98.39 130.8 66.57 101.8 203.4 65.67
-DCStereo* [34] 8.38 2.77 1.54 1.82 1.19 3.99 1.09 2.64 2.09 1.04 2.65
-LightEndoStereo(Ours) 7.62 2.316 1.669 2.281 1.025 4.428 1.167 2.809 2.125 0.4793 2.592
-TABLE II: Results of comparison experiments on the SERV-CT dataset.Methods marked with (*) reference results from other papers. Bold font indicates the optimal results.
-Model MAE(pixel)$\downarrow$ D1$\downarrow$ Bad1$\downarrow$ Bad2$\downarrow$ Bad3$\downarrow$
-GwcNet [16] 5.813 33.92% 79.37% 56.32% 40.93%
-PSMNet* [29] 6.355 - 70.88% 49.80% 33.87%
-RAFT [21] 21.500 87.50% 95.80% 91.80% 88.50%
-IGEV [20] 32.895 92.91% 98.28% 96.53% 94.76%
-DLNR [19] 22.030 88.97% 96.21% 92.61% 89.13%
-Selective-Stereo [31] 24.103 86.36% 95.68% 91.33% 87.01%
-MSDESIS-resnet32 [7] 5.742 28.26% 78.81% 55.44% 37.79%
-MSDESIS-light [7] 7.065 41.35% 81.09% 63.17% 49.22%
-LightEndoStereo(Ours) 2.367 14.34% 58.59% 31.52% 18.49%
-We evaluated our algorithm against several methods for natural image stereo matching and two algorithms for endoscopic data.
-Specifically, GwcNet [16], PSMNet [29], RAFT [35], IGEV [20], DLNR [19], and Selective-Stereo [31] are methods designed for natural images, with DLNR and Selective-Stereo focusing on optimizing high-frequency details in disparity maps.
-For these natural image methods, except for PSMNet, we retrained them on the SCARED dataset.
-For PSMNet, we directly cited the results from [34, 3].
-MSDESIS [7] and DCStereo [34] are algorithms tailored for endoscopic images.
-DCStereo’s evaluation methodology on SCARED aligns with ours, so we directly cited their results.
-However, since DCStereo was not tested on the SERV-CT dataset and its code is not publicly available, we omitted their results on SERV-CT.
-MSDESIS was trained on the SCARED dataset, and we directly loaded their pre-trained weights.
-Given that our maximum disparity is set to 192, whereas MSDESIS uses a maximum disparity of 320, the results we reproduced differ slightly from those reported in their paper.
-The evaluation results on SCARED are shown in Table I, and the results on SERV-CT are shown in Table II.
-LightEndoStereo achieved SOTA performance on both datasets and demonstrated excellent generalization on SERV-CT.
-IV-D Runtime Evaluation
-TABLE III: Quantitative results of model computational performance.
-Note that only methods with an average MAE < 3mm on the SCARED dataset are included.Methods marked with (*) reference results from other papers.
-Bold font indicates the optimal results.
+GwcNet 9.07 2.92 1.43 1.70 1.20 3.86 1.13 2.93 1.99 0.61 2.68
+PSMNet* 8.96 2.77 1.43 1.83 1.42 3.99 1.08 2.82 1.95 0.56 2.68
+RAFT 7.92 2.36 1.72 2.16 1.88 4.36 1.08 2.87 1.98 1.54 2.79
+IGEV 7.95 2.39 1.69 2.26 1.93 3.89 1.17 2.89 2.50 1.25 2.79
+DLNR 7.84 2.52 2.65 3.20 3.41 5.37 1.91 4.32 4.82 10.25 4.63
+Selective-Stereo 7.53 2.40 1.66 2.31 2.26 3.86 0.95 2.83 1.77 1.01 2.66
+MSDESIS-full 10.14 3.46 3.30 3.88 2.09 5.42 1.63 3.37 1.76 0.63 3.57
+MSDESIS-light 11.45 15.91 7.72 10.67 9.84 98.39 130.80 66.57 101.80 203.40 65.67
+DCStereo* 8.38 2.77 1.54 1.82 1.19 3.99 1.09 2.64 2.09 1.04 2.65
+RRESM(ours) 7.62 2.31 1.66 2.28 1.02 4.42 1.16 2.80 2.12 0.47 2.59
+In contrast, MSDESIS [20] and DCStereo [13] are methods tailored for endoscopic stereo matching.
+For DCStereo, we directly cited its results on the SCARED dataset, as their evaluation protocol is consistent with ours.
+However, since DCStereo was not evaluated on the SERV-CT dataset and its implementation is not publicly available, we omitted it from SERV-CT comparisons.
+As for MSDESIS, we used the pre-trained model provided by the authors.
+Note that our method uses a maximum disparity of 192, while MSDESIS adopts 320, which may account for slight differences between our reproduced results and those reported in their original paper.
+Figure 5: Visualization of disparity estimation on the SCARED and SERV-CT datasets.
+(a) and (b) are from SCARED, while (c) and (d) are from SERV-CT.
+In cases with limited depth variation, such as (b), most methods perform similarly.
+However, in high-frequency regions like surgical tool-tissue boundaries in (a), RRESM yields more accurate depth predictions.
+On SERV-CT, our method also delivers competitive results.
+Note: MAE is measured in millimeters ($m$) for SCARED (with ground-truth depth) and in pixels for SERV-CT (with ground-truth disparity).
+The quantitative results on the SCARED and SERV-CT datasets are presented in Table 2 and Table 1, respectively.
+Sample visualizations are shown in Fig.
+5.
+On the SCARED dataset, RRESM achieves the best average MAE of 2.59 mm, outperforming both natural image and endoscopy-specific baselines.
+On the SERV-CT dataset, our method also achieves superior performance across all metrics, including the lowest MAE (2.367 pixels), lowest Bad-n error rates, and a significantly lower D1 score.
+These results highlight RRESM’s robustness in handling challenging tissue boundaries and its strong generalization ability to unseen clinical data.
+4.4 Runtime Evaluation
+Table 3: Quantitative results of model computational performance.
 Model Params(M) FLOPs(T) Runtimes(ms)
-GwcNet [16] 6.43 2.466 26.4
-PSMNet* [29] 3.672 - 225
-RAFT [21] 11.11 3.757 468.6
-IGEV [20] 12.5 3.288 434.5
-Selective-Stereo [31] 13.141 4.062 516.52
-DCStereo* [34] 3.404 - 191
-LightEndoStereo(Ours) 11.094 0.846 23.38
-To evaluate the runtime performance of our model, we conducted tests using a single Nvidia RTX 2080Ti GPU, paired with an Intel(R) Xeon(R) Gold 6130T CPU @ 2.10GHz.
-We sampled 100 images ($1280\times 1024$) from dataset 8 of the SCARED dataset for testing and used the average inference time as the evaluation metric.
-The results are shown in Table III.
+GwcNet 6.43 2.466 26.4
+PSMNet* 3.672 - 225
+RAFT 11.11 3.757 468.6
+IGEV 12.5 3.288 434.5
+Selective-Stereo 13.141 4.062 516.52
+DCStereo* 3.404 - 191
+RRESM(ours) 11.094 0.846 23.38
+To evaluate the runtime performance of our model, we conducted tests using a single Nvidia RTX 2080Ti GPU.
+We sampled 100 images ($1280\times 1024$) from dataset 8 of the SCARED for testing and used the average inference time as the evaluation metric.
+The results are shown in Table 3.
 Although the MSDESIS-Light model has the smallest footprint, its matching accuracy is poor.
-In contrast, LightEndoStereo achieved a real-time matching speed of 23.38 ms per frame.
+In contrast, RRESM achieved a real-time matching speed of 23.38 ms per frame.
+4.5 Ablation Study
+Table 4: Ablation Study Results. The average MAE on the SCARED dataset is used as the accuracy evaluation metric.
+MCA HFDO MAE($m$) Runtime(ms) Params(M) FLOPs(T)
+2.718 25.412ms 13.047 1.755
+$\checkmark$ 2.645 25.972ms 13.077 1.757
+2.621 22.655ms 11.064 0.843
+$\checkmark$ 2.592 23.386ms 11.094 0.846
+We conducted ablation studies to evaluate the contributions of the proposed 3D Mamba Coordinate Attention (MCA) and High-Frequency Disparity Optimization (HFDO) modules.
+MCA is designed to improve cost aggregation efficiency by modeling axis-specific long-range dependencies, while HFDO focuses on refining depth estimates in regions with high-frequency disparity transitions, such as tissue-tool boundaries.
 Figure 6: Visualization of disparity optimization in high-frequency regions. The top row shows the depth maps, while the bottom row displays the corresponding error maps.
-TABLE IV: Ablation Study Results.
-We use the average MAE on the SCARED dataset as the accuracy evaluation metric.
-When ablating the 3D UNet-MCA, we employ the widely-used 3D Hourglass aggregation network.
-3D UNet-MCA HFDO MAE(mm) Runtime(ms) Params(M) FLOPs(T)
-2.718 25.412ms 13.0479 1.7552
-$\checkmark$ 2.645 25.972ms 13.0773 1.7579
-2.621 22.655ms 11.0649 0.8435
-$\checkmark$ 2.592 23.386ms 11.0943 0.8461
-IV-E Ablation Study
-We conducted detailed ablation studies on the proposed 3D MCA and HFDO modules (see Table IV).
-When ablating the 3D UNet-MCA aggregation network, we employ the widely-used 3D Hourglass aggregation network [16].
-The 3D Hourglass aggregation network is a stacked symmetric multi-scale network that features a large number of parameters and high computational complexity.
-As shown in Table III, the 3D UNet-MCA not only reduces the number of parameters and computational complexity but also enhances disparity estimation accuracy.
-As Figure 4 illustrates, the HFDO module significantly reduces matching errors at tissue boundaries, leading to notable improvements in both the Stacked 3D Hourglass and 3D UNet-MCA configurations.
-V Conclusions and limitations
-We introduce LightEndoStereo, a novel lightweight real-time stereo matching algorithm tailored for endoscopic images.
-Our approach integrates a coordinate attention mechanism and the Mamba block to achieve efficient cost aggregation via a lightweight 3D UNet.
-Moreover, by enhancing high-frequency features in the wavelet domain, we optimize disparity estimation, effectively mitigating large matching errors at tissue boundaries—a common challenge in endoscopic stereo matching.
-During validation, LightEndoStereo exhibited strong generalization on the SERV-CT dataset, a potential that warrants further exploration.
-Acknowledging the significance of generalizable stereo matching algorithms, we intend to delve deeper into this area in our future research.
+The experiments were performed on the SCARED dataset using our 3D U-Net-based architecture.
+To assess MCA, we replaced it with a conventional stacked 3D Hourglass network [10], which is commonly used in existing stereo matching pipelines.
+For evaluating HFDO, we simply removed the wavelet-based refinement module from the full pipeline.
+We report the Mean Absolute Error (MAE) in millimeters, runtime per frame, model parameters, and FLOPs, as summarized in Table 4.
+Results show that both modules individually contribute to improved accuracy.
+Incorporating MCA alone reduces MAE from 2.718 to 2.621 mm while significantly lowering model complexity.
+Adding HFDO further improves performance to 2.592 mm.
+Notably, the complete model achieves the best trade-off between accuracy and efficiency.
+As visualized in Fig.
+6, HFDO particularly enhances boundary precision by reducing local matching errors in high-frequency regions.
+These results validate the effectiveness and complementary nature of MCA and HFDO within the overall architecture.
+5 Discussion and Conclusions
+We present RRESM, a real-time stereo matching framework tailored for endoscopic imagery.
+By integrating a 3D Coordinate Attention mechanism with the Mamba block, RRESM enables efficient and lightweight cost aggregation while capturing long-range dependencies.
+The High-Frequency Disparity Optimization module further enhances disparity accuracy near anatomical edges through wavelet-based refinement.
+Extensive experiments on the SCARED and SERV-CT datasets show that RRESM achieves state-of-the-art performance and generalizes well to unseen domains, running at 42 FPS on high-resolution inputs.
+Despite these promising results, our method has several limitations.
+Its performance may degrade under extreme illumination variation, specular reflections, or endoscope lens occlusion.
+Additionally, the current pipeline is limited to stereo input and assumes spatial rectification.
+In future work, we will (1) explore uncertainty-aware disparity refinement, (2) extend RRESM to multi-view and monocular depth estimation, and (3) validate its robustness across diverse anatomical sites and surgical conditions.